OpenStudy (anonymous):

Raul plans to spend no more than $78.00 on two shirts and a pair of jeans. He bought the two shirts for $19.89 each. How much can he spend on the jeans can someone please help me

OpenStudy (triciaal):

shirt =s jeans = j 2s + 2 j less or = 78; 1s = 19.89 2s = 2 x 19.89 = 2j < or = 78 - 2s

OpenStudy (anonymous):

so 2 shirts = 39.78

OpenStudy (texaschic101):

2s + j <= 78 (less then or equal ) 2(19.89) + j <= 78 39.78 + j <= 78 j <= 78 - 39.78 j <= 38.22 Raul can spend $ 38.22 or less on jeans

OpenStudy (anonymous):

Thank you so much texaschic101

OpenStudy (triciaal):

thanks yes it would be better to put just J not 2j and only < 78 not = and < 78 nevertheless, the amount for jeans < $38.22

OpenStudy (texaschic101):

glad I could help :)